given utility function U= where PX = $12, $, py = $4 and the income of the consumer is, M = $240.
A, Find the utility maximizing combinations of X and Y.
B, calculate marginal rate of substitution of X for Y (MRSX,Y) at equilibrium and interpret your result.
1
Expert's answer
2020-06-10T19:02:46-0400

Since the utility function should be a function of two variables, then


"MU_x=\frac {\delta U}{\delta x}""MU_y=\frac {\delta U}{\delta y}"


"12x+4y=240"


"\frac {MU_x}{12}=\frac {MU_y}{4}"


"MU_x=3MU_y"